a Corporation issued $600,000, 796, 10-year E10-8 . On August 1, 2014, Orteg Instructions Prepare journal entries to record the following events. bonds r face value. Interest is payable annually on August 1. Ortegas year-end is December 31at (a) The issuance of the bonds. (b) The accrual of interest on December 31, 2014. (c) The payment of interest on August 1, 2015
Answer #1

Date

Accounts title and explanation

Debit

Credit

Aug 1 2014..

Cash

$ 6,00,000.00

           Bonds payable

$ 6,00,000.00

(To record issue of bonds)

Dec 31 2014..

Interest on bonds

$      17,500.00

    Interest payable

$      17,500.00

(To record interest on bonds payable)

Aug 1 2015..

Interest payable

$      17,500.00

Interest on bonds

$      24,500.00

      Cash

$      42,000.00

(To record payment of interest on bond)

Accrued interest on bonds will be recorded for 5 months starting from august till December.

Working

Date

Accounts title and explanation

Debit

Credit

Aug 1 2014..

Cash

600000

           Bonds payable

=+BJ103

(To record issue of bonds)

Dec 31 2014..

Interest on bonds

=(600000*7%)/12*5

    Interest payable

=(600000*7%)/12*5

(To record interest on bonds payable)

Aug 1 2015..

Interest payable

=(600000*7%)/12*5

Interest on bonds

=(600000*7%)/12*7

      Cash

+17500+24500

(To record payment of interest on bond)